# Gaussian process modelling of infectious diseases using the Greta software package and GPUs

### v1

@@ 3.1 Data
The UK government releases weekly reports of notifiable infectious diseases (NOIDS) (29) in England and Wales. Several diseases are reported: Measles, Mumps, Rubella, Scarlet Fever, Whooping Cough, Tuberculosis, Acute Meningitis, Malaria, and food poisoning. The reports contain how many cases there were of each disease within Public Health England (PHE) Region; county; and local and unitary authorities. National park local authorities and some corporations in London were not included in the weekly reports. We chose to focus on Tuberculosis counts in the local and unitary authorities in the East and West Midland PHE regions between 2022 and 2024, which had a significant number of counts to model and could demonstrate how Gaussian Process models can be used on current outbreaks. Note that the cases only include active Tuberculosis cases and not latent Tuberculosis cases. For training the Gaussian Process model we used the years 2022 and 2023, and for testing the predictive ability used the first month of 2024. This means we had 104 weeks of training data and 4 weeks of testing data. We also took data from the government on the longitude and latitude (and boundaries) (30) and the population size (31) of each local planning authority.
@@ 3.2 Likelihood
The data consists of Tuberculosis counts in authority $i$ in week $j$ for $j\in{1,\ldots,104}$ in the East and West Midlands, England. We model these counts using the Negative Binomial distribution, which has the following likelihood:
@@ 3.2.1 Zero-inflated likelihood
3.3 Relationship to the Gaussian Process
In larger populations, we would expect a higher absolute number of infections. As a result, it is common to split the mean number of infections ($\mu_{ij}$) as a product of the background population effect ($e_{ij}$) and the excess risk ($\theta_{ij}$) (32) as to focus on modelling the excess risk:
$\mu_{ij}=e_{ij}\theta_{ij}.$
$e_{ij}=p_{ij}R$
$R=\frac{\sum_{i}\sum_{j}y_{ij}}{\sum_{i}\sum_{j}p_{ij}}$
$\log(\theta_{ij})=f(x_{ij})$
$f(.)\sim GP(m(.),k(.,.)).$
We take the mean function to be a constant that we estimate from the data, while the kernel function is specified as described in the next section.
@@ 3.4 Priors
Part of the Bayesian framework is setting prior distribution on the model parameters. For the kernel parameters we use the priors recommended by the Stan handbook (33), such that:
@@
also suggested by Stan in their Developer Wiki (34) which makes $\phi$ relatively small and consequently we have an overdispersed model. We also gave the mean function parameter the prior in (3.4).
For the zero-inflated models we have an extra parameter $\pi$ that gives the probability of extra zeroes. For this, we used the recommendation of (35), which suggested defining $\pi$ as a function of $\lambda$
@@ 3.5 Models
Rather than building every combination of models from the set of kernel functions, models are altered sequentially based on the best previous model. The focus is initially on optimising the temporal kernel through smoothness and adding a periodic kernel function, while keeping the spatial kernel fixed. Then the smoothness of the spatial kernel can be varied to further optimise. Finally we test how changing the distribution to Zero-Inflated affects the model. The exact choices of kernel functions can be seen in Table 1.
@@ 5 $\text{mat}_{32}$ + periodic $\text{mat}_{32}$ $k_{time}*k_{space}$ Zero-Inflated Negative Binomial
A grid approach was used to defining the inducing points, because it is computationally cheap and most of the variation in our explanatory variables comes from the time component as the locations are the same throughout the data. We kept all the locations and sampled the time component every 5 weeks, with the final week also included as it is the most important week for prediction. So in total we used 22 time points, giving 1430 (65 locations * 22 weeks) inducing inputs.
@@ 3.7 Results
Model comparison and assessment was conducted using the LOOIC, Tukey statistic and CPRS over 26 weeks of prediction (Table 4). CPRS across weeks is shown in Figure 1. There was relatively small differences across models, with Model 2 overall giving the best trade-off between model complexity and fit. Full summaries of the six models are presented in the Supplementary materials.
@@
Figure 2: Map showing observed (shaded map) and median posterior predictions (shaded circles) of Tuberculosis cases for the East and West Midlands for the first week of 2024 using our best model (model 2). This map was made using Leaflet (36) with Shiny (37).
@@ 4 Conclusions
Much previous work has been done to introduce computationally efficient and accurate approximations to these inversions. For example Vecchia approximation (38); H-matrices (39, 40) amongst others. As well as efficient linear algebra, the use of efficient computational hardware to speed up model fitting and prediction can drastically improve the practical fitting of GPs, and software enabling this with minimal user input and expertise is important.

### v3

@@ 3.1 Data
The UK government releases weekly reports of notifiable infectious diseases (NOIDS) (GOV.UK, 2023) in England and Wales. Several diseases are reported: Measles, Mumps, Rubella, Scarlet Fever, Whooping Cough, Tuberculosis, Acute Meningitis, Malaria, and food poisoning. The reports contain how many cases there were of each disease within Public Health England (PHE) Region; county; and local and unitary authorities. National park local authorities and some corporations in London were not included in the weekly reports. We chose to focus on Tuberculosis counts in the local and unitary authorities in the 65 East and West Midland PHE regions between 2022 and 2024, which had a significant number of counts to model and could demonstrate how Gaussian Process models can be used on current outbreaks. This gave a total of 6760 training data points. Note that the cases only include active Tuberculosis cases and not latent Tuberculosis cases. For training the Gaussian Process model we used the years 2022 and 2023, and for testing the predictive ability used the first month of 2024. This means we had 104 weeks of training data and 4 weeks of testing data. We also took data from the government on the longitude and latitude (and boundaries) (Office for National Statistics (2022a), ONS) and the population size (Office for National Statistics (2022b), ONS) of each local planning authority.
@@ 3.2 Likelihood
The data consists of Tuberculosis counts in authority $i$ for $i\in{1,\ldots,65}$ in week $j$ for $j\in{1,\ldots,104}$ in the East and West Midlands, England. We model these counts using the Negative Binomial distribution, which has the following likelihood:
@@ 3.3 Priors
Part of the Bayesian framework is setting prior distribution on the model parameters. For the kernel parameters we use the priors recommended by the Stan handbook (Stan Development Team, 2024), such that:
@@
also suggested by Stan in their Developer Wiki (Stan Development Team, 2023) which makes $\phi$ relatively small and consequently we have an overdispersed model. We also gave the mean function parameter the prior in (3.3).
For the zero-inflated models we have an extra parameter $\pi$ that gives the probability of extra zeroes. For this, we used the recommendation of R-INLA Project (2024), which suggested defining $\pi$ as a function of $\lambda$
@@ 3.4 Models
Rather than building every combination of models from the set of kernel functions, models are altered sequentially based on the best previous model. The focus is initially on optimising the temporal kernel through varying smoothness and also allowing a periodic kernel function, while keeping the spatial kernel fixed (Models 1-3). The Matèrn(3,2) kernel is one with a mid-level smoothness as a starting point and often the standard kernel used in spatial statistics. Once the optimal temporal structure of those testes is found, namely a Matèrn(3,2) added to a periodic kernel, then different spatial kernels are also tested (Models 4&6). The smoothness of the spatial kernel, based on Euclidian distance between the latitude and longitude of each authority, can be varied to further optimise model fit. Finally we test how changing the distribution to Zero-Inflated affects the model (Model 5). The exact choices of kernel functions can be seen in Table 1.
@@ 5 $\text{mat}_{32}$ + periodic $\text{mat}_{32}$ $k_{time}*k_{space}$ Zero-Inflated Negative Binomial
A grid approach was used to defining the inducing points, because it is computationally cheap and most of the variation in our explanatory variables comes from the time component as the locations are the same throughout the data. We kept all the locations and sampled the time component every 5 weeks, with the final week also included as it is the most important week for prediction. So in total we used 22 time points, giving 1430 (65 locations * 22 weeks) inducing inputs. The HMC algorithm was run for 2000 iterations, with the first 1000 discarded as burn-in. Since we used a Bayesian method we needed to check convergence of the chains to a stable distribution, running four independent chains. Brooks-Gelman-Rubin statistics were monitored until less than 1.1, supporting convergence (Brooks and Gelman, 1998) (see Supporting Information for example traceplots and the convergence statistics).
@@ 3.6 Results
Model comparison and assessment was initially conducted using the LOOIC, Tukey statistic and CPRS over 4 weeks of prediction (see Supporting Information), but was extended to up to 26 weeks ahead (Table 4). CPRS across weeks is shown in Figure 2. There was relatively small differences across models, with Model 2 overall giving the best trade-off between model complexity and fit. Full summaries of the six models are presented in the Supplementary materials.
@@
Figure 3: Map showing observed (shaded map) and median posterior predictions (shaded circles) of Tuberculosis cases for the East and West Midlands for the first week of 2024 using our best model (model 2). This map was made using Leaflet (Graul, 2016) with Shiny (Chang et al., 2024).
@@ 4 Conclusions
Much previous work has been done to introduce computationally efficient and accurate approximations to these inversions. For example Vecchia approximation (Vecchia, 2018); H-matrices (Geoga et al., 2020; Moran and Wheeler, 2022) amongst others. As well as efficient linear algebra, the use of efficient computational hardware to speed up model fitting and prediction can drastically improve the practical fitting of GPs, and software enabling this with minimal user input and expertise is important.
@@
The approach was capable of predicting at least 26 weeks ahead without significant loss of predictive ability, suggesting the training data are sufficient to learn signal through space and time. The support for a periodic signal of tuberculosis spread in the East and West Midlands suggests seasonal variation in incidence that is relatively consistent across years, supporting previous work on TB globally Fares (2011).